OpenStudy (anonymous):

what is an equation where the answer is 26?

OpenStudy (anonymous):

x + 6 = 20 does this work?

OpenStudy (hoblos):

x-6=20 **

OpenStudy (anonymous):

a three step equation

OpenStudy (hoblos):

three step equation?? what does this mean

OpenStudy (anonymous):

an equation where three step s are needed to solve the equation

OpenStudy (anonymous):

a thankless task i assure you

OpenStudy (anonymous):

hoblos, it's something from the very early days of algebra. I actually had to look it up because I'd forgotten haha, but it's basically something involving three steps to the form of: simply and collect like terms; rearrange for x; solve for x.

OpenStudy (anonymous):

yea so do u have an equation for me?

OpenStudy (anonymous):

Oh and yeah, it's like scrubbing toilets. Nobody will thank you or appreciate the time you spend doing it.

OpenStudy (anonymous):

i will please give me an equation

OpenStudy (anonymous):

try \[2x+1=x+27\]

OpenStudy (anonymous):

or maybe \[3x+1=x+53\]

OpenStudy (anonymous):

\[\Large 87,613x - 1,914,068 - 13,995x = 0\]

OpenStudy (anonymous):

no satellite they both equal 13 not 26

OpenStudy (anonymous):

fewscrews those number sare too big give smaller numbers

OpenStudy (anonymous):

And both satellite's answers are 26.

OpenStudy (anonymous):

jeez my elementary algebra must have deserted me. lets try and solve \[3x+1=x+53\] 1) subtract x from both sides get \[2x+1=53\]2) subtract 1 from both sides get \[2x=52\]3) divide by 2 get \[x=26\] hmmm

OpenStudy (anonymous):

noooo

OpenStudy (anonymous):

lilnick32, do you understand changing sides, collecting like terms, etc?

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

yea he is right

OpenStudy (anonymous):

Ok. Say you have the equation:\[\Large 3x - 30 = 0\]You want to move the 30 to the other side, since it doesn't involve the variable or co-efficient x. So you add 30 to both sides:\[\Large 3x - 30 + 30 = 0 + 30\]But of course most people don't write it out like that, and simply jump to \[\Large 3x = 30\]Now you know that 3 times x is equal to 30, so to find x, divide 30 by 3:\[\Large x = \frac{{30}}{3}\]And you have your answer:\[\Large x = 10\]

OpenStudy (anonymous):

now this is the last question im asking. so what is an equation in which the answer is 28?

OpenStudy (anonymous):

Read everything everyone has posted here and try and find one =)

OpenStudy (anonymous):

please this is the last one

OpenStudy (anonymous):

Fine. But just remember, when you actually need to use this kind of thing, you're not going to be able to simply ask somebody for help.\[\Large 2x + 13 = x + 41\]

OpenStudy (anonymous):

thank u